OpenStudy (anonymous):

Last year, Carlos had $10,000 to invest. He invested some of it in an account that paid 5% simple interest per year, and he invested the rest in an account that paid 6% simple interest per year. After one year, he received a total of %540 in interest. How much did he invest in each account?

OpenStudy (anonymous):

$540 in interest?

OpenStudy (anonymous):

yes sorry i had a typo

OpenStudy (anonymous):

Let x represent the amount invested at 5% simple interest I=Prt, t=1, so I=Pr 540=.05x+.06(10000-x)

OpenStudy (anonymous):

need more help?

OpenStudy (anonymous):

no i think i got it and then after i plug in x in y=10000-x? so i can find y?

OpenStudy (anonymous):

When you get x, you have the amount invested @ 5%. (10000-x) is the amount invested @ 6%

OpenStudy (anonymous):

now i am lost

OpenStudy (anonymous):

Ok, we'll solve for x

OpenStudy (anonymous):

540=.05x+.06(10000-x)

OpenStudy (anonymous):

.05x+600-.06x=540 -.01x=-60 x=-60/-0.01=$6000 invested at 5%

OpenStudy (anonymous):

Now the amount invested at 6% is (10000-6000)=$4000

OpenStudy (anonymous):

To check your answer, plug in 6000 for x in the original: .05(6000)+.06(4000)=

OpenStudy (anonymous):

oh okay thank you! would you be able to help me with a similiar question ??

OpenStudy (anonymous):

x can represent anything you want it to. we can also use it to represent the amount invested @ 6%, but have to rewrite the equation to this: 540=.06x+.05(10000-x) Now x will be the amount invested at 6%, and (10000-x) will be the amount invested @ 5%

OpenStudy (anonymous):

Chang bought a desktop computer and a laptop computer. Before finance charges, the laptop cost $300 more than the desktop. He paid for the computers using two different financing plans. For the desktop the interest rate was 9% per year, and for the laptop it was 7% per year. The total finance charges for one year were $365 . How much did each computer cost before finance charges?